# KGGen: Extracting Knowledge Graphs from Plain Text with Language Models

### v1

@@ 5 Results
We use MINE to benchmark KGGen against leading existing methods of plain-text-to-KG extraction: OpenIE Angeli et al. (2015) and GraphRAG Larson & Truitt (2024). After providing this quantitative comparison of extraction fidelity, we present qualitative results demonstrating the advantages of KGGen over past methods.
5.1 Evaluations on MINE
Figure 3 displays accuracies from KGGen, OpenIE, and GraphRAG on MINE. Figure 4 shows an example query from MINE and relevant relations extracted by KGGen, OpenIE, and GraphRAG.
Figure 3: Distribution of MINE scores across 100 articles for GraphRAG, OpenIE, and KGGen. Dotted vertical lines show average performance. KGGen scored $66.07\%$ on average, significantly outperforming GraphRag $47.80\%$ and OpenIE $29.84\%$.
Figure 4: An example query from the MINE benchmark, along with relevant relations in the KGs extracted by KGGen, GraphRAG, and OpenIE. Note that the relation triples extracted by KGGen contain the fact being queried for, whereas the KGs extracted by GraphRAG and OpenIE do not. The relation types extracted by KGGen are more concise and generalize more easily than those from GraphRAG and OpenIE. The full article that these relations were extracted from can be found in Appendix D.
@@ 5.2 Qualitative Results
As seen in Figure 5(b) and 5(e), GraphRAG often generates a minimal number of nodes and connections for an entire article. This sparsity results in the omission of critical relationships and information. For compression, Figure 5(a) and 5(d) illustrate sections of the KGs generated by KGGen for the same articles. Figure 5(c) illustrates one of many issues in OpenIE’s KGs. Firstly, most nodes are unreasonably long, incoherent phrases. Many of these nodes are redundant copies of one another, adding unnecessary complexity to the graph. Additionally, as seen in 5(f) OpenIE frequently produces generic nodes such as “it” and “are.” Due to their frequency, these nodes, which contain no useful information, often end up as some of the most well-connected nodes in the graph. By contrast, KGGen consistently generates KGs that are dense and coherent, effectively capturing critical relationships and information from the articles.
(a) Section of KG generated by KGGen on “How Music Influences Mood”
(b) Full KG generated by GraphRAG on “How Music Influences Mood”
(c) Section of KG generated by OpenIE, on “How Music Influences Mood”, with most node labels omitted for readability.
(d) Section of KG generated by KGGen on “The Chemistry of Cooking”
(e) Full KG generated by GraphRAG on “The Chemistry of Cooking”
(f) Section of KG generated by OpenIE on “The Chemistry of Cooking”
@@
6 Future Work
We propose MINE – the first benchmark for KG extraction from plain text. To solve the data-shortage hindering development of graph-based foundation models, we present KGGen, a plain-text-to-KG extractor that outperforms existing approaches by up to $18\%$ on MINE.
Although KGGen beats existing methods by significant margins, the graphs still exhibit problems, like over or under-clustering. More research into better forms of clustering could improve the quality of our KGs. Additionally, our benchmark, MINE, currently measures performance on relatively short corpora, whereas KGs are primarily used to handle massive amounts of information efficiently. Future expansions of our benchmark could focus on larger corpora to better measure the practicality of different extraction techniques.
7 Related Work
Interest in automated methods to produce structured text to store ontologies dates back to at least 2001 when large volumes of plain text began to flood the fledgling internet (Maedche & Staab, 2001). KG extraction from unstructured text has seen significant advances through rule-based and LM-powered approaches in the last 15 years. Early work (Suchanek et al., 2007) used hard-coded rules to develop YAGO, a KG extracted from Wikipedia containing over five million facts, and rules-based extraction still has appeal for those producing KGs in multi-modal domains today (Norabid & Fauzi, 2022; Oramas et al., 2015). With the development of modern natural language processing, hard-coded rules generally ceded to more advanced approaches based on neural networks. For instance, OpenIE (Angeli et al., 2015) provides a two-tiered extraction system: first, self-contained clauses are identified by a classifier; then, Angeli et al. run natural logic inference to extract the most representative entities and relations from the identified clauses. Stanford KBP (Angeli et al., 2013) presents another seminal early approach to using deep networks for entity extraction.
As early as 2015, some hypothesized that extracting KGs would go hand-in-hand with developing better language models (Domeniconi et al., 2015). More recently, evidence has emerged that transformer-based architectures can identify complex relationships between entities, leading to a wave of transformer-based KG extraction techniques, which range from fully automatic (Qiao et al., 2022; Arsenyan et al., 2023; Zhang & Soh, 2024) to human-assisted (Kommineni et al., 2024). Our contribution to the extraction literature is to build KGs conducive to embedding algorithms such as TransE and TransR (Bordes et al., 2013; Lin et al., 2015). We observed that when one extracts KGs from plaintext, the nodes and relations are often so specific that they are unique. This causes the estimation of embeddings to be under-specified. We develop a method for automatic KG extraction from plain text that clusters similar nodes and edges to prevent this under-specification. This leads to a KG with better connectivity and more functional nodes and edges.
Evaluating the quality of knowledge graphs is important to ensure usefulness and reliability in downstream applications. Early evaluation methods focused primarily on directly assessing aspects such as completeness and connectivity or using rule-based statistical methods, while recent approaches emphasize usability in downstream applications and incorporation of semantic coherence(Xue & Zou, 2023).
In the late 2000s, research focused on assessing the correctness and consistency of KGs. The evaluations relied on expert annotations by selecting random facts from the generated KG and then calculating the accuracy of those facts. (Suchanek et al., 2007) This proved to be laborious and prone to errors. This led to accuracy approximation methods like KGEval (Ojha & Talukdar, 2017) and Two-State Weight Clustering Sampling(TWCS) (Gao et al., 2018), which employed sampling methods with statistical guarantees as well as use less annotation labor. As the KGs became larger and more diverse, particularly with the rise of automated extraction techniques from web data, this generated more pressure on annotators, leading to methods like Monte-Carlo search being used for the interactive annotation of triples (Qi et al., 2022). Furthermore, because accuracy alone did not fully capture the complexity of the knowledge graph, more evaluation metrics like completeness were used to characterize the quality of knowledge graphs. (Issa et al., 2021).
Two notable approaches to KG evaluation are the LP-Measure and the triple trustworthiness measurement (KGTtm) model. LP-Measure assesses tDhe quality of a KG through link prediction tasks, eliminating the need for human labor or a gold standard (Zhu et al., 2023). This method evaluates KGs based on their consistency and redundancy by removing a portion of the graph and testing whether the removed triples can be recovered through link prediction tools. Empirical evidence suggests that LP-Measure can effectively distinguish between “good” and “bad” KGs. The KGTtm model, on the other hand, evaluates the coherence of triples within a knowledge graph Jia et al. (2019). Based on these evaluation methods, frameworks like Knowledge Graph Evaluation via Downstream Tasks(KGrEaT) and DiffQ(differential testing) emerged. KGrEaT provides a comprehensive assessment of KGs by evaluating their performance on downstream tasks such as classification, clustering, and recommendation (Heist et al., 2023) rather than focusing solely on correctness or completeness. In contrast, DiffQ uses embedding models to evaluate the KG’s quality and assign a DiffQ Score, resulting in improved KG quality assessment. Tan et al. (2024)
This shift towards task-based evaluation underscores the importance of usability and accessibility in KGs. Factors such as expressiveness, context information, and ease of integration into downstream AI applications are now central to evaluating their quality and effectiveness.
8 Acknowledgments
@@
JK acknowledges support from NSF grant number DGE-1656518. SK acknowledges support from NSF 2046795 and 2205329, the MacArthur Foundation, Stanford HAI, and Google Inc.
@@ Appendix A Prompts for KG Extraction
The initial KG is extracted using the following two prompts.
Prompt for extracting entities: Extract key entities from the given text. Extracted entities are nouns, verbs, or adjectives, particularly regarding sentiment. This is for an extraction task, please be thorough and accurate to the reference text. Prompt for extracting relations: Extract subject-predicate-object triples from the assistant message. A predicate (1-3 words) defines the relationship between the subject and object. Relationship may be fact or sentiment based on assistant’s message. Subject and object are entities. Entities provided are from the assistant message and prior conversation history, though you may not need all of them. This is for an extraction task, please be thorough, accurate, and faithful to the reference text.
After extracting the entities and relations from each unit of text, we begin the clustering process, which is performed using the following prompts.
Prompt for clustering entities:
Find ONE cluster of related entities from this list. A cluster should contain entities that are the same in meaning, with different:
- tenses
- plural forms
- stem forms
- upper/lower cases
Or entities with close semantic meanings.
Return only if you find entities that clearly belong together.
If you can’t find a clear cluster, return an empty list.
Prompt for validating node clusters: Verify if these entities belong in the same cluster.
A cluster should contain entities that are the same in meaning, with different:
- tenses
- plural forms
- stem forms
- upper/lower cases
Or entities with close semantic meanings.
Return the entities that you are confident belong together as a single cluster.
If you’re not confident, return an empty list.
Prompt for clustering edges
Find ONE cluster of closely related predicates from this list.
A cluster should contain predicates that are the same in meaning, with different:
- tenses
- plural forms
- stem forms
- upper/lower cases
Predicates are the relations between subject and object entities. Ensure that the predicates in the same cluster have very close semantic meanings to describe the relation between the same subject and object entities.
Return only if you find predicates that clearly belong together.
If you can’t find a clear cluster, return an empty list.
Prompt for validating cluster edges
Verify if these predicates belong in the same cluster.
A cluster should contain predicates that are the same in meaning, with different:
- tenses
- plural forms
- stem forms
- upper/lower cases
Predicates are the relations between subject and object entities. Ensure that the predicates in the same cluster have very close semantic meanings to describe the relation between the same subject and object entities.
Return the predicates that you are confident belong together as a single cluster.
If you’re not confident, return an empty list.
Appendix B Validation of KG Extraction
This section provides the LLM generations used to validate our KG extraction method.
Prompt for extracting entities: Extract key entities from the given text. Extracted entities are nouns, verbs, or adjectives, particularly regarding sentiment. This is for an extraction task, please be thorough and accurate to the reference text. Prompt for extracting relations: Extract subject-predicate-object triples from the assistant message. A predicate (1-3 words) defines the relationship between the subject and object. Relationship may be fact or sentiment based on assistant’s message. Subject and object are entities. Entities provided are from the assistant message and prior conversation history, though you may not need all of them. This is for an extraction task, please be thorough, accurate, and faithful to the reference text.
Appendix C Prompts for MINE
@@
Appendix D Example Article from MINE

### v2

@@ 6 Results
We use MINE to benchmark KGGen against leading existing methods of plain-text-to-KG extraction: OpenIE Angeli et al. (2015) and GraphRAG Larson and Truitt (2024). After providing this quantitative comparison of extraction fidelity, we present qualitative results that demonstrate the advantages of KGGen over past methods.
Figure 3: Distribution of MINE-1 scores across 100 articles for GraphRAG, OpenIE, and KGGen. Dotted vertical lines show average performance. KGGen scored $66.07\%$ on average, significantly outperforming GraphRag $47.80\%$ and OpenIE $29.84\%$.
Figure 4: An example query from the MINE-1 benchmark, along with relevant relations in the KGs extracted by KGGen, GraphRAG, and OpenIE. Note that the relation triples extracted by KGGen contain the fact being queried for, whereas the KGs extracted by GraphRAG and OpenIE do not. The relation types extracted by KGGen are more concise and generalize more easily than those from GraphRAG and OpenIE. The full article that these relations were extracted from can be found in Appendix C.
6.1 Evaluations on MINE-1
Figure 4 displays accuracies from KGGen, OpenIE, and GraphRAG on MINE. Note that KGGen outperforms competing methods. Figure 4 shows an example query from MINE-1 and relevant relations extracted by KGGen, OpenIE, and GraphRAG.
6.1.1 Generalization Across Language Models
To evaluate KGGen’s robustness across different foundation models, we tested its performance on MINE-1 using multiple state-of-the-art LLMs. Table 1 shows that KGGen maintains strong performance across different models, with Claude Sonnet 3.5 achieving the highest score of 73%.
Table 1: Performance comparison of KGGen
(a) KGGen performance on MINE-1 across different language models
Model MINE-1 Score (%)
Claude Sonnet 3.5 73
GPT-4o 66
Gemini 2.0 Flash 44
(b) Validity of extracted triples across different methods
Method Valid Triples (%)
KGGen 98/100 (98%)
GraphRAG 0/100 (0%)
OpenIE 55/100 (55%)
KGGen’s extraction methodology generalizes well across different foundation models. Although Claude Sonnet 3.5 achieves the highest score of 73%, all tested models maintain reasonable extraction quality, making KGGen adaptable to users’ preferred LLM providers.
6.2 Evaluations on MINE-2: RAG performance
Figure 5 shows comparable performance between KGGen and GraphRag on MINE-2.
Figure 5: Comparison between KGGen and GraphRag on MINE-2. The two methods perform comparably.
6.3 Evaluation on Human-Annotated Data: SemEval-2010
To evaluate KGGen’s extraction quality against human-annotated ground truth, we conducted experiments using the SemEval-2010 Task 8 dataset. We randomly selected 100 sentences from the dataset, each containing two manually labeled target entities. After removing entity markup tags to ensure unbiased extraction, we applied KGGen to extract entities and relationships from each sentence.
Our evaluation focused on entity extraction accuracy, as the dataset’s relation labels consist of only 5 broad categorical types rather than specific semantic relations. We assessed whether both human-annotated target entities appeared in KGGen’s extracted entities, allowing for more specific entity descriptions (e.g., “Eurasia exhibition” instead of just “exhibition”) that still referred to the same object.
Results show that KGGen successfully captured both target entities in 96% of cases (96/100). The method consistently extracted more detailed entity descriptions compared to human annotations, often identifying additional relevant entities in longer sentences. For example, given the sentence “The ambitious Eurasia exhibition arose from an idea by Achille Bonito Oliva,” with target entities ‘exhibition’ and ‘idea’, KGGen extracted [‘Eurasia exhibition’, ‘idea’, ‘Achille Bonito Oliva’], providing more specific and complete entity identification.
@@ 6.4 Qualitative Results
We first evaluated the fundamental quality of extracted knowledge graphs by examining whether the extracted triples conform to the basic definition of a knowledge graph: subject-predicate-object triples where subjects and objects are entities (nodes) and predicates are relationships (edges). We randomly selected 100 triples from each method and manually evaluated their validity. The results are shown in Table 1.
Despite GraphRAG’s comparable performance on downstream tasks, it does not extract structures that closely resemble traditional knowledge graphs, which is a major strength of KGGen. As seen in Figures 6(b) and 6(e), GraphRAG often extracts very few nodes and connections for an entire article. This sparsity results in the omission of critical relationships and information. For compression, Figure 6(a) and 6(d) illustrate sections of the KGs generated by KGGen for the same articles. Figure 6(c) illustrates one of many issues in OpenIE’s KGs. Firstly, most node types are hyperspecific, incoherent phrases. Many of these nodes are redundant near-copies of each other, adding unnecessary complexity to the graph. Additionally, as seen in 6(f) OpenIE primarily uses pattern matching to identify entities, and frequently produces generic nodes such as “it" and “are". Due to their frequency, these nodes, which contain no useful information, often end up as some of the most well-connected nodes in the graph. Consequently, unrelated concepts end up being just two hops apart, linked by paths through nodes like “it" or “are". By contrast, KGGen consistently generates KGs that are informative and coherent, effectively capturing critical relationships and information from the articles.
6.5 A Note on Scaling
A major motivation for the creation of KGGen was to produce graphs where edge types are generalizable, and used more than once when the corpus grows large. To test the re-usability of our relations, we generate three knowledge graphs from text of different sizes: 10 000 characters, 100 000 characters, and 1 000 000 characters and plot the number of edges divided by the number of unique relations. As one can see from Figure 7, KGGen reuses each relation-type an average of $10$ times, and the average number of occurrences of each relation-type increases with the size of the corpus. By contrast, GraphRAG reuses each relation type an average of $2$ times regardless of the size of the graph. This suggests that the relations in GraphRAG do not generalize as the corpus grows.
(a) Section of KG generated by KGGen on “How Music Influences Mood"
(b) Full KG generated by GraphRAG on “How Music Influences Mood"
(c) Section of KG generated by OpenIE, on “How Music Influences Mood", with most node labels omitted for readability.
(d) Section of KG generated by KGGen on “The Chemistry of Cooking"
(e) Full KG generated by GraphRAG on “The Chemistry of Cooking"
(f) Section of KG generated by OpenIE on “The Chemistry of Cooking"
@@ 6.5 A Note on Scaling
6.6 Efficiency and Cost Analysis
To evaluate the practical applicability of KGGen, we analyzed its computational efficiency and cost on a large-scale extraction task. We extracted a KG from the novel The Name of the Wind by Patrick Rothfuss, processing text corpora of increasing sizes. Table 2 demonstrates KGGen’s de-duplication effectiveness across different scales.
Table 2: KGGen scaling characteristics showing entity and relation de-duplication
Corpus Size Pre- Post- Entity Pre- Post- Relation Pre- Post- Edge
(chars) Entities Entities De-dup Ratio Relations Relations De-dup Ratio Edges Edges De-dup Ratio
100 1 1 1.000 1 1 1.000 1 1 1.000
1,000 20 18 0.900 12 12 1.000 9 9 1.000
10,000 90 78 0.867 78 75 0.962 62 57 0.919
100,000 727 604 0.831 926 924 0.998 498 424 0.851
1,000,000 4,602 3,573 0.776 8,137 8,094 0.995 3,180 2,448 0.770
The de-duplication ratios improve with scale, demonstrating the effectiveness of our clustering algorithm. For the complete novel (1M characters), KGGen achieved a 22.4% reduction in entities and 23% reduction in edges through intelligent consolidation. This analysis highlights the contribution of each module: the extraction phase (Steps 1-2) captures comprehensive information from text, while the resolution phase (Step 3) significantly reduces redundancy without information loss. Table 3 presents the computational cost breakdown for processing the entire novel:
Table 3: KGGen cost and throughput analysis for 1M character corpus
Step Prompt Completion Total Time Throughput Cost
Tokens Tokens Tokens (s) (tokens/s) (\$)
KG Extraction (Steps 1–2) 1.59M 0.63M 2.22M 273 8,139 0.46
Entity/Edge Resolution (Step 3) 2.93M 0.22M 3.15M 279 11,304 0.38
Total 4.52M 0.85M 5.37M 551 9,739 0.84
For comparison, we evaluated GraphRAG on the same corpus: results can be found in Table 4.
Table 4: GraphRAG scaling characteristics on the same corpus
Corpus Size (chars) Entities Relations Edge Types Time (s)
100 2 1 1 1.89
1,000 4 3 3 3.01
10,000 16 20 20 29.71
100,000 80 100 99 205.12
1,000,000 514 981 966 2,079.17
While GraphRAG is faster on short corpora, its execution time scales superlinearly. The complete extraction time comparison reveals a significant difference: GraphRAG requires 2,319 seconds for the extraction phase alone on the 1M character corpus, compared to KGGen’s total processing time of 551 seconds (including both extraction and resolution). Additionally, GraphRAG produces nearly as many relation types (966) as edges (981), indicating minimal relation reuse and poor generalization compared to KGGen’s efficient consolidation.
7 Broader Impact and Community Adoption
Our work produces a KG-from-plain-text extractor that helps to solve the KG-scarcity problem. Improved knowledge-graph extractors could lead to the prevalence of structured text, which can help improve factuality and reliability of information retrieval systems. Our open-source implementation has already enjoyed widespread community adoption. The package has received over 700 Github stars and has been downloaded over 12,000 times since its release.
8 Limitations and Future Work
Although KGGen holds many advantages over past extraction methods, its graphs still exhibit problems, like over or under de-duplication of entities and relations. Further research into entity resolution could improve the quality of our KGs. Additionally, our benchmarks measure corpora of up to 5M tokens, which does not reflect the size of web-scale text that would be necessary to produce a KG foundation model. Future expansions of our benchmark could focus on larger corpora to better measure the practicality of different extraction techniques.
Domain-specific knowledge extraction presents additional challenges. Fields like medicine and finance require specialized domain knowledge that general-purpose LLMs may lack, potentially limiting extraction quality compared to human experts. While MINE-2 demonstrates KGGen’s capability across diverse domains, incorporating domain-specific ontologies could improve extraction precision. Future work could explore adaptive ontology integration to balance structure with completeness.
Acknowledgments and Disclosure of Funding
@@ 8 Limitations and Future Work
We acknowledge support from NSF grant numbers DGE-1656518, NSF 2046795 and 2205329, the MacArthur Foundation, Stanford HAI, and Google Incorporated.
@@ Appendix A Prompts for KG Extraction
The initial KG was extracted using the following two prompts passed as DSPy signature descriptions.
Prompt for extracting entities: Extract key entities from the source text. Extracted entities are subjects or objects. This is for an extraction task, please be thorough and accurate to the reference text. Prompt for extracting relations: Extract subject-predicate-object triples from the source text. Subject and object must be from entities list. Entities provided were previously extracted from the same source text. This is for an extraction task, please be thorough, accurate, and faithful to the reference text.
After extracting the entities and relations from each unit of text, we begin the de-duplication process, which is performed using the following prompts.
Prompt for entity or edge resolution:
Find duplicate {item_type} for the item and an alias that best represents the duplicates. Duplicates are those that are the same in meaning, such as with variation in tense, plural form, stem form, case, abbreviation, shorthand. Return an empty list if there are none.
Appendix B Prompts for MINE
@@
Prompt for RAG response: Use the following knowledge graph triples and text evidence to answer the question.
Triples: {triples_text}
Text Evidence: {text_block}
Question: {query} Answer:
Prompt for Evaluating containment of WikiQA answer: You are a fact-checking assistant
Question: question
Expected answer: expected
Model’s response: response
Does the model’s response contain the information in the expected answer?
Respond with one word: Yes or No.
Appendix C Example Article from MINE